OpenStudy (anonymous):

Find G if...... limx→−2 (x^2+(G−1)x+G)/(x2) =1

OpenStudy (anonymous):

x+2 on the bottom not x2

OpenStudy (freckles):

G is a constant? And that the difference of G and 1 in that one group of parenthesis right?

OpenStudy (freckles):

\[\lim_{x \rightarrow -2}\frac{x^2+(G-1)x+G}{x+2}=1\]

OpenStudy (freckles):

I think something maybe wrong with this question because if this limit were to exist you should have 0/0 and if you have 0/0 you should also be able to use l'hosptal rule and I'm getting two different G's

OpenStudy (anonymous):

Thats the equations and i need to find G, but I'm not sure where to start

OpenStudy (freckles):

well you should have 0/0 when pluggin -2

OpenStudy (freckles):

start there

OpenStudy (anonymous):

the denominator would be 0 so i cant solve for G

OpenStudy (freckles):

For \[\lim_{x \rightarrow -2}\frac{x^2+(G-1)x+G}{x+2} \text{ \to exist } \] you must have 0/0 when pluggin in -2

OpenStudy (freckles):

so \[(-2)^2+(G-1)(-2)+G=0 \]

OpenStudy (anonymous):

what about the bottom (x+2), what do i do with that?

OpenStudy (amistre64):

does x+2 = 0 when x=-2?

OpenStudy (anonymous):

Yes

OpenStudy (freckles):

And also l'hopstail will work just differentiated wrong :p

OpenStudy (amistre64):

then we dont need to do anything with it at the moment ...

OpenStudy (anonymous):

ok...

OpenStudy (amistre64):

\[\frac{(-2)^2+(G-1)(2)+G~=~0}{(-2)+2~=~0}\] solve for G

OpenStudy (amistre64):

ugh, dropped a -

OpenStudy (freckles):

\[L=\lim_{x \rightarrow a}\frac{f(x)}{g(x)}\] If L is a number we do have to have f(a)=0 and g(a)=0 or f(a)=some constant and g(a)=some constant not equal to 0

OpenStudy (freckles):

obviously we have the first case since g(a) does equal 0 so we have to find when f(a)=0

OpenStudy (freckles):

Though f(a)=0 and g(a)=0 might give other conclusions than L being a number

OpenStudy (freckles):

but here we are assuming it will give us an L that exists since L=1

OpenStudy (amistre64):

4 - 2G +2 + G = 0 6 - G = 0

OpenStudy (freckles):

Try @amistre64 's answer there and you will see it will give you the limit is 1

OpenStudy (anonymous):

but im confused of the -2+2=0 and what we do with it since its on the bottom causing an error because we cant divide by 0

OpenStudy (amistre64):

it only causes an error of x+2 is NOT a factor of the toip

OpenStudy (amistre64):

*top

OpenStudy (freckles):

We are basically trying to chose a G such that we have a hole at x=-2 and not a vertical asymptote

OpenStudy (amistre64):

if (x+2) is a common factor, we can essentially cancel it out with algebra

OpenStudy (freckles):

to get a hole we have to have 0/0 and not c/0 where c is not equal to 0

OpenStudy (amistre64):

\[\frac{(x+2)~f(x)}{(x+2)}\] when x=-2, top and bottom go to 0

OpenStudy (freckles):

we want to write \[\frac{x^2+(G-1)x+G}{x+2}=\frac{(x+r)(x+2)}{(x+2)}=x+r \]

OpenStudy (freckles):

oops beat me to it

OpenStudy (anonymous):

ok, so G=-6?

OpenStudy (freckles):

yes

OpenStudy (freckles):

wait no

OpenStudy (freckles):

6

OpenStudy (amistre64):

it was half right :)

OpenStudy (anonymous):

oh i see where i went wrong, so G=6 and that is my answer?

OpenStudy (freckles):

you want to see another way

OpenStudy (anonymous):

yes please, if you dont mind

OpenStudy (freckles):

\[\frac{x^2+(G-1)x+G}{x+2}=\frac{(x+r)(x+2)}{(x+2)}=x+r \] So when is \[x^2+(G-1)x+G=(x+r)(x+2)\] \[x^2+(G-1)x+G=x^2+(r+2)x+2r\] so G-1=r+2 and G=2r replace G with 2r in first equation 2r-1=r+2 2r-r=1+2 r=3 so G=2r=2(3)=6

OpenStudy (anonymous):

ok thank you both very much!!

OpenStudy (freckles):

There is also another way if you know l'hospital

OpenStudy (freckles):

\[2x+(G-1)=1 \text{ as } x \text{ approaches } -2 \\2(-2)+(G-1)=1 \\-4+G-1=1 \\G-5=1 \\ G=6 \]

OpenStudy (freckles):

@amistre64 check out my pretty system of linear equations way :P

OpenStudy (anonymous):

ok, i have not learned that yet but it helps to see the answer

OpenStudy (amistre64):

i had concidered Lhopping first, but you had already started with the long way ;)

OpenStudy (freckles):

the long way is pretty too

OpenStudy (freckles):

tkhunny suggest using long division

OpenStudy (freckles):

you could have done that too

OpenStudy (freckles):

or synthetic division also

OpenStudy (amistre64):

tkhunny is never to be trusted .... never ever ever!

OpenStudy (freckles):

which is basically the first way

OpenStudy (amistre64):

@tkhunny i spose its fair if you make it a tag :)

OpenStudy (freckles):

x ------------------------------- x+2| x^2+(G-1)x+G -(x^2+2x) --------------- (G-1-2)x+G so then x+(G-3) ------------------------------- x+2 | x^2+(G-1)x+G -(x^2+2x) ------------------ (G-3)x+G -((G-3)x+2(G-3)) ---------------- G-2(G-3) That remainder is G-2(G-3) =G-2G+6 =-G+6 And the remainder is suppose to be 0 so -G+6=0 when G=6

OpenStudy (tkhunny):

I may not have understood the question. Are you sure you typed it the same the second time? This is why we like @amistre64

OpenStudy (amistre64):

im devious ... lol

OpenStudy (anonymous):

hey, @freckles and @amistre64 where did you get the 'r' from??

OpenStudy (freckles):

\[x^2+bx+c=(x+r_1)(x+r_2)\]

OpenStudy (freckles):

where r1 and r2 are the zeros of x^2+bx+c

OpenStudy (freckles):

just like if we had a cubic \[x^3+bx^2+cx+d=(x+r_1)(x+r_2)(x+r_3)\]

OpenStudy (freckles):

where r is a complex number since not all quadratics are factorable over the reals

OpenStudy (freckles):

we knew we wanted \[x^2+(G-1)x+G \text{ to factor as } (x+2)(x+r)\] so we could cancel the x+2 on bottom we just called it r to begin with because we didn't know what the other zero needed to be at the time

OpenStudy (anonymous):

ok thanks!!